OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 14 and a13 = –58?

OpenStudy (cwrw238):

a13 = 14 +12d -58 = 14 + 12d d = -72 / 12 = -6 32nd term = 14 + (-6*31) = 14 - 186 = -172